### v3

@@ 5.1 Shape of the dark matter halo
Figure 8: The DM halo shape, $q_{\rm DM}=Z/X$ versus $p_{\rm DM}=Y/X$, of the Milky Way in comparison with TNG50 and Auriga galaxies. The red dashed contours indicate $1\sigma$ and $3\sigma$ confidence level of the Milky Way DM axis ratios obtained by our models, the red star marks the best-fitting model with $q_{\rm DM}=0.92$ and $p_{\rm DM}=0.8$. Each black dot is a Milky Way-like galaxy from TNG50 simulations, each black multiple is a galaxy from Auriga simulations. The plus at the bottom left indicate the typical uncertainty of DM halo shapes measured for TNG50/Auriga galaxies. The four simulations highlighted by red/magenta circles/multiple are those associated a vertical satellite plane as we will shown later, those highlighted by black circles are all simulations lay within $3\sigma$ range of the Milky Way but without significant vertical satellite planes.
@@
Figure 9: The DM axis ratio measured at three different radii, $r=15,30,50$ kpc, for TNG50 and Auriga galaxies, in comparison with that of the Milky Way obtained by our model allowing $p_{\rm DM}(r)$ and $q_{\rm DM}(r)$ vary as a function of radius. This model has large uncertainty and we only weakly constrain the DM axis ratios at different radii, which shows a trend to be more oblate in the inner regions and to be more vertically aligned in the outer regions, consistent with the trend shown in the Milky Way analogues, TNG50 501208, TNG50 546474, Auriga 12, and TNG50 483294. For these four galaxies, the arrows point to their intrinsic axis ratios measured at the principle axes of the DM halo. These four galaxies have twist DM halo and being vertically aligned at $r\gtrsim 20$ kpc like we obtained for the Milky Way, bur their intrinsic DM shape are common comparing to the general TNG50/Auriga galaxies if we hypothetically tilt their disk and inner DM halo to the long-intermediate axes plane of the outer DM halo.
@@
We take 198 Milky Way-like or M31-like galaxies (Pillepich2024MNRAS.535.1721P) from the TNG50 simulations (Nelson2019ComAC...6....2N; Nelson2019MNRAS.490.3234N; Pillepich2019MNRAS.490.3196P) and 30 galaxies from Auriga (Grand2017MNRAS.467..179G), and we perform further selections to enhance their comparability with the Milky Way. We calculate the flattening of the stellar component $q_{\rm star}=z_{\rm gc}/x_{\rm gc}$ of each simulation and define the disk fraction by considering stars with circularity $\lambda_{z}>0.5$ as part of the disk. Note here we do not define a dynamically cold disk with $\lambda_{z}>0.7$ (Genel2015ApJ...804L..40G) or $\lambda_{z}>0.8$ (Zhu2018NatAs...2..233Z), instead, we include some stars on dynamically warm orbits to make the disk more comparable to the Milky Way disk which was defined by fitting exponential profiles to the stellar number density distribution (Grand2017MNRAS.467..179G). The Milky Way has $f_{\rm disk}\sim 0.7$ and global $q_{\rm star}\sim 0.15$ (Bland-Hawthorn2016). We selected galaxies using a slightly loose threshold with $f_{\rm disk}>0.6$ and $q_{\rm star}<0.3$ to retain a relatively large sample. This selection removes galaxies that have experienced relatively recent major mergers, which create a more prominent bulge or stellar halo than that of the Milky Way. We further removed a few galaxies with ongoing major mergers that have not stabilised their stellar components. This results in 94 TNG50 galaxies and 11 Auriga galaxies.
@@
We calculate the intrinsic shape of the DM halo in the coordinate $\boldsymbol{X}\equiv\{X,Y,Z\}$, defined exactly the same way as we defined it for the Milky Way in our model. We first define the $Z$ axis as vertical to the stellar disk. Then, we project the DM halo along the $Z$ axis and define the long and short axes in the disk plane as $X$ and $Y$, respectively. We measure the intrinsic shape of the DM halo in the fixed coordinate system and measure the axis ratios $p_{\rm DM}=Y/X$ and $q_{\rm DM}=Z/X$ at $r=15,30,50$ kpc, respectively. In this definition, $p_{\rm DM}$ is restricted to be less than unity, while $q_{\rm DM}$ is allowed to be either lower or higher.
@@
We further show the DM axis ratios of TNG50/Auriga galaxies measured at different radii in Fig. 9 and compare them with the Milky Way DM axis ratio obtained by the model with varying $p_{\rm DM}(r)$ and $q_{\rm DM}(r)$. The above three galaxies, with average axis ratios within $1\sigma$, actually have twisted DM halos that are more oblate in the inner regions and become more vertically aligned in the outer regions. TNG50 483594, outside of $1\sigma$ but within $3\sigma$, exhibits a similar trend but is intrinsically rounder. Although not statistically significant, the Milky Way DM halo is likely to show a similar trend, as suggested by our model with varying $p_{\rm DM}(r)$ and $q_{\rm DM}(r)$. The radial variation of the DM halo shape we found, although tentative, is consistent with recent results from the modelling of the Sagittarius stream (Vasiliev2021MNRAS.501.2279V).
@@ 5.2 Arrangement of the satellite system
Figure 10: Orbit configuration of satellite galaxies for the Milky Way and the four analogues highlighted in Fig.8 and Fig.9. We show orbits of the ten brightest or most massive satellites in YZYZ planes. The blue disk at top left indicates the orientation of stellar disk. The analogues exhibit striking similarities with the Milky Way that DM halo coincides with a plane of satellites and vertically aligns with the Galactic disk.
@@
The so-called ‘satellite plane’ of the Milky Way is especially notable in the orbits of the most massive satellites (Sawala2023NatAs...7..481S). We first illustrate the orbital configuration of the massive satellite galaxies in the Milky Way, as well as those in the 13 TNG50/Auriga galaxies whose DM shapes fall within the $3\sigma$ confidence level of the Milky Way.
@@
For the Milky Way, we selected the ten brightest satellites and used their observed 3D positions and velocities as initial conditions, as per Pace2022ApJ...940..136P, to compute their orbital paths within our derived optimal gravitational potential. For each of the TNG50/Auriga galaxies, we take the ten most massive satellites and use their actual trajectories back approximately one orbital period from $z=0$ in most cases. For the satellites that have been orbited for many periods and are already very close to the galaxy centre at $z=0$, we took an orbital period from 3 Gyr ago for better visualisation.
@@
In Fig.10, we show these four galaxies that have a nearly oblate outer DM halo vertical to the stellar disk: TNG50 501208, TNG50 546474, Auriga 12, and TNG50 483594. They exhibit vertical satellite planes similar to the Milky Way; the first three are the only galaxies with DM shapes that fall within the $1\sigma$ confidence level of the Milky Way. There are no clear satellite planes in the remaining nine galaxies within $3\sigma$ (see Appendix A).
@@
In general, a correlation is evident between the DM halo shape and the arrangement of the satellite systems in TNG50 galaxies (see Appendix A). Most TNG50 galaxies have oblate DM halos with long-intermediate axes aligned with the stellar disk and paired with coplanar satellite systems. While galaxies with almost spherical DM halos display satellites distributed in a nearly isotropic fashion, those with DM halos aligned vertically to the stellar disk tend to have satellite systems organised vertically as well. This correlation is expected as a result of the DM halo assembling over time through the accretion of satellite galaxies with similar net angular momentum, consistent with previous findings from other cosmological simulations (Shao2016MNRAS.460.3772S).
@@
5.3 A vertically orientated dark matter halo in alignment with a satellite plane results from a flip of the stellar disk
@@
Figure 11: A spiral-in minor merger flips the stellar disk illustrated by TNG50 501208. The left column show orbits of the ten most massive mergers colored their maximum history mass. The red dashed with $M_{\rm max}=5.3\times 10^{10}$M  is a major merger finished at about 8.5 Gyr ago, the red solid with $M_{\rm max}=7.1\times 10^{10}$M  is the most massive minor merger followed. The rest columns show the DM distribution (the gray shadow), the projection of the stellar disk (the blue ellipse with size enlarged five times for visualization) and the orbits progress of the most massive minor merger (red solid curve) at different snapshots in the history, the axis ratios of the DM halo measured at $r=50$ kpc are labeled. A thin stellar disk was formed right after the major merger finished at about 8.5 Gyr ago. At about t=67t=6-7 Gyr ago the disk was aligned almost co-planer with the long-intermediate axis plane of the DM halo and orbital plane of the satellites. The disk was flipped in the last seven Gyr, associated with the spiral-in minor merger (red solid curve), while the general shape of DM halo (illustrated by $q_{\rm DM}$ and $p_{\rm DM}$ measured at 50 kpc in a fixed coordinate) and satellite arrangement remain unchanged, thus resulting the vertically alignment between the stellar disk and DM halo/satellite orbit plane at present (also see Movie S1).
@@
Figure 12: The evolution history of TNG50 501208 (left) and TNG50 546474 (right). We describe the four panels with TNG50 501208. Top: assembly history of the galaxy, the solid thick curves represent two of the most massive mergers the galaxy experienced, the dashed curves represent small mergers. TNG50 501208 has assembled $80\%$ of its DM mass and $72\%$ of its stellar mass at $t=6$ Gyr ago. Second row: Distance of the satellite to the galacetic center as a function of time. The black solid curve is a major merger ended at $t\sim 8.5$ Gyr ago, the red curve is the following most massive minor merger with $M_{\rm max}=7.1\times 10^{10}$ M, the vertical red dashed line indicate the time it reached the pericenter for the first time. Third row: The variation of disk spin direction as a function of time. The disk was co-aligned with the long-intermediate axis plane of the DM halo at $t\sim 6$ Gyr ago, and it has tilted $80^{\circ}$ in the past 6 Gyr. Bottom: Evolution of the DM axis ratios, with $p_{\rm DM}=Y/X$ and $q_{\rm DM}=Z/X$, the DM shape remain largely unchanged in the past 8 Gyr. The four panels in the right show similar evolution history of TNG50 546474.
@@
We further trace the evolutionary history of the four Auriga/TNG50 galaxies, TNG50 501208, TNG50 546474, TNG50 483594, and Auriga 12, and find that the distinct vertical alignment of the DM halo/satellite plane is due to the tilt of the stellar disk with a large angle, which we call a “flip of the disk”.
@@
To check the evolutionary history, we pay special attention to the merger history, the evolution of the stellar disk, and the shape of the DM halo. We calculate the spin direction of all stars at $r<20$ kpc at each snapshot in the history and define $\theta_{\rm disk}$ as the angle between the spin vector and its projection line in the XY plane. In this definition, we have $\theta_{\rm disk,z=0}=90^{\circ}$. We then calculate the DM axis ratios at 50 kpc ($p_{\rm DM}=Y/X$, $q_{\rm DM}=Z/X$) in different snapshots. $XYZ$ is the coordinate defined at $z=0$, with the $Z$ axis perpendicular to the disk, as we defined previously.
@@
As illustrated in Fig.11 and quantified in Fig. 12, the last major merger of TNG50 501208 ended $t\sim 8.5$ Gyr ago. A stellar disk was formed gradually after the major merger. At $t\sim 6-7$ Gyr ago, the stellar disk was almost aligned in the long-intermediate axis plane of the DM halo; if not perturbed by other mergers, the system would stabilize in this configuration. The galaxy was followed by a series of minor mergers, the most massive one having $M_{\rm max}=7.1\times 10^{10}$ Mcoming in on an orbit that is retrograde to the disk rotation. At $t=6$ Gyr ago, when the massive minor merger reached its pericenter for the first time, the disk started to tilt, initiating continuous tilting in the subsequent six Gyr. The disk has since flipped about $80^{\circ}$ until $z=0$, with an average rate of $\sim 13^{\circ}$ per Gyr, while it was not significantly perturbed. Although the satellite perturbed the DM halo’s shape at pericenter, the DM halo quickly re-stabilized. The mass accreted from this satellite and other minor mergers continuously built the DM halo along the orbit, consistent with the pre-existing DM halo shape; the axis ratios of the DM halo measured at $r=50$ kpc remained largely unchanged. The disk flip resulted in the DM halo and stellar disk’s vertical alignment at $z=0$.
@@
The satellite galaxy that spiralled inward brings a substantial net angular momentum. It is important to note that a tilt rate of $\sim 13^{\circ}$ per Gyr corresponds to about 2 km/s at a radius of $r=10$ kpc, indicating a very minor rigid motion of the disk. This scenario requires torques that transfer less than $\lesssim 1\%$ of the angular momentum of the satellite galaxy, which has a total mass comparable to that of the stellar disk. In our analysis of 13 TNG/Auriga galaxies within a $3\sigma$ threshold, eight galaxies exhibited significant disk tilts, with tilt angles ranging from $30^{\circ}$ to $100^{\circ}$. Seven of them are likely caused by minor mergers with $M_{\rm max}>10^{10}\,M_{\odot}$, mostly spiralled in (see appendix A). Disk flips are fairly common in cosmological simulations (Dillamore2022MNRAS.513.1867D; Earp2017MNRAS.469.4095E; Bett2012MNRAS.420.3324B). Direct N-body simulations show that the torque on the stellar disk from a satellite galaxy is a promising mechanism that induces the disk tilt (Dodge2023MNRAS.518.2870D). Previous analysis on Auriga galaxies show disk tile is correlated with time of close satellite interactions (Gomez2017MNRAS.472.3722G), consistent with our results. It is also found that even low mass satellites that penetrate the outer regions of a galaxy can significantly perturb and tilt a host galactic disk (Gomez2017MNRAS.465.3446G), due to not only direct tidal perturbation but also to the generation of asymmetric features in the DM halo that can be efficiently transmitted to its inner regions.
@@
In a similar pattern, the evolution of TNG50 546474 is presented in the right panel of Fig. 12. This galaxy underwent a major merger that ended around $t\sim 7.5$ Gyr ago, immediately followed by the most massive minor merger with $M_{\rm max}=1.4\times 10^{10}$M. At $t\sim 7$ Gyr ago, after the reformation of the disk post-major merger and stabilisation of the DM halo, the disk and DM halo were nearly coplanar and aligned. Subsequently, the DM halo’s shape saw minimal change, although it gained about $30\%$ of its mass in the upcoming 7 Gyr through minor mergers on orbits consistent with the pre-existing DM halo shape. At the same time, the disk has tilted approximately $60^{\circ}$ with an average rate of $\sim 8^{\circ}$ per Gyr.
@@
Similar figures for Auriga 12 and TNG50 483594 are shown in Figure 13. Auriga 12 has had more frequent recent mergers. Its stellar disk was coplanar aligned with the DM halo at $t\sim 8-10$ Gyr ago; the torques from the combination of a few subsequent minor mergers have tilted the disk by $\sim 90^{\circ}$. The DM halo was influenced but has maintained a similar shape since about 10 Gyr ago.
@@
The DM shape of TNG50 483594 is outside the $1\sigma$ but within the $3\sigma$ confidence level of the Milky Way. This galaxy is special in that it was quickly assembled with many head-on mergers at $t\gtrsim 8$ Gyr ago, but without a GSE-like major merger, which resulted in a rather spherical DM halo. Due to frequent mergers in the early universe, it has quickly assembled stellar mass and halted star formation. With a near-spherical DM halo, its stellar disk was arbitrarily orientated at t=610t=6-10 Gyr ago. The subsequent minor mergers exhibit a satellite plane whose contribution has not significantly changed the DM halo shape until $t\sim 6$ Gyr ago, along with a massive minor merger with $M_{\rm max}=4.5\times 10^{10}$M. The minor mergers built the halo along their orbits and tilted the disk by about $30^{\circ}$ in another direction, leading to the vertical alignment of the DM halo and the disk at $z=0$. In this case, the tilt of the disk is not the only reason; it strengthens the misalignment and contributes to the vertical alignment of the stellar disk and the satellite distribution observed at $z=0$. Even without the disk tilt, there is still a misalignment of the stellar disk and the satellite plane due to variations in the filament network along with mass accreted into the galaxy, as shown in (Shao2021).
@@
Despite the tilting of the disk, the DM halo of the inner regions ($r\lesssim 20$ kpc) was tilting with the stellar disk. As a result, the DM halos of these galaxies tend to be twisted, being oblate in the inner regions, and more vertically aligned with the disk in the outer regions, as already shown in Figure 9, consistent with the trend we obtained for the Milky Way.
@@
The three $1\sigma$ galaxies, TNG50 501208, TNG50 546474, and Auriga 12 would transform into a typical galaxy with an oblate DM halo by hypothetically tilting their stellar disk (and inner DM halo at $r\lesssim 20$ kpc) back to the long-intermediate axis plane of the DM halo in the outer regions. Note that TNG50 403894 also has a twisted DM halo; however, its transition radius, similar to other cases of twisted halos caused by variations of filament accretion (Shao2021), is larger than the three cases caused purely by the tilt of the disk (see Fig.9).

### v4

@@ 5.1 Shape of the dark matter halo
Figure 8: DM halo shape, $q_{\rm DM}=Z/X$ versus $p_{\rm DM}=Y/X$, of the Milky Way in comparison with TNG50 and Auriga galaxies. The dashed red contours indicate the $1\sigma$ and $3\sigma$ confidence level of the Milky Way DM axis ratios obtained by our models. The red star marks the best-fitting model with $q_{\rm DM}=0.92$ and $p_{\rm DM}=0.8$. Each black dot is a Milky Way-like galaxy from TNG50 simulations. Each black multiple is a galaxy from Auriga simulations. The plus at the bottom left indicates the typical uncertainty of DM halo shapes measured for TNG50/Auriga galaxies. The four simulations highlighted by red/magenta circles/multiple are those associated with a vertical satellite plane. As we show later, those highlighted by black circles are all simulations lying within a $3\sigma$ range of the Milky Way but without significant vertical satellite planes.
@@
Figure 9: DM axis ratio measured at three different radii, $r=15,30,50$ kpc, for TNG50 and Auriga galaxies, in comparison with that of the Milky Way obtained by our model allowing $p_{\rm DM}(r)$ and $q_{\rm DM}(r)$ vary as a function of radius. This model has a large uncertainty and we only weakly constrain the DM axis ratios at different radii, which shows a trend to be more oblate in the inner regions and to be more vertically aligned in the outer regions, consistent with the trend shown in the Milky Way analogues, TNG50 501208, TNG50 546474, Auriga 12, and TNG50 483294. For these four galaxies, the arrows point to their intrinsic axis ratios measured at the principle axes of the DM halo. These four galaxies have twisted DM halos and are vertically aligned at $r\gtrsim 20$ kpc, as we obtained for the Milky Way, bur their intrinsic DM shapes are common compared to the general TNG50/Auriga galaxies if we hypothetically tilt their disc and inner DM halo to the long-intermediate axes plane of the outer DM halo.
@@
We took 198 Milky Way-like or M31-like galaxies (Pillepich et al., 2024) from the TNG50 simulations (Nelson et al., 2019b, a; Pillepich et al., 2019) and 30 galaxies from Auriga (Grand et al., 2017), and we performed further selections to enhance their comparability with the Milky Way. We calculated the flattening of the stellar component, $q_{\rm star}=z_{\rm gc}/x_{\rm gc}$, of each simulation and defined the disc fraction by considering stars with circularity $\lambda_{z}>0.5$ as part of the disc. Note here that we did not define a dynamically cold disc with $\lambda_{z}>0.7$ (Genel et al., 2015) or $\lambda_{z}>0.8$ (Zhu et al., 2018). Instead, we included some stars on dynamically warm orbits to make the disc comparable to the Milky Way disc that was defined by fitting exponential profiles to the stellar number density distribution (Grand et al., 2017). The Milky Way has $f_{\rm disc}\sim 0.7$ and global $q_{\rm star}\sim 0.15$ (Bland-Hawthorn and Gerhard, 2016). We selected galaxies using a slightly loose threshold with $f_{\rm disc}>0.6$ and $q_{\rm star}<0.3$ to retain a relatively large sample. This selection removed galaxies that have experienced relatively recent major mergers, which create a more prominent bulge or stellar halo than that of the Milky Way. We further removed a few galaxies with ongoing major mergers that have not stabilised their stellar components. This results in 94 TNG50 galaxies and 11 Auriga galaxies.
@@
We calculated the intrinsic shape of the DM halo in the co-ordinate $\boldsymbol{X}\equiv\{X,Y,Z\}$, defined exactly the same way as we defined it for the Milky Way in our model. We first defined the $Z$ axis as being vertical to the stellar disc. Then, we projected the DM halo along the $Z$ axis and defined the long and short axes in the disc plane as $X$ and $Y$, respectively. We measured the intrinsic shape of the DM halo in the fixed co-ordinate system and measured the axis ratios $p_{\rm DM}=Y/X$ and $q_{\rm DM}=Z/X$ at $r=15,30,50$ kpc, respectively. In this definition, $p_{\rm DM}$ is restricted to be less than unity, while $q_{\rm DM}$ is allowed to be either lower or higher.
@@
We further show the DM axis ratios of TNG50/Auriga galaxies measured at different radii in Fig. 9 and compare them with the Milky Way DM axis ratio obtained by the model with varying $p_{\rm DM}(r)$ and $q_{\rm DM}(r)$. The above three galaxies, with average axis ratios within $1\sigma$, actually have twisted DM halos that are more oblate in the inner regions and become more vertically aligned in the outer regions. TNG50 483594, outside of $1\sigma$ but within $3\sigma$, exhibits a similar trend but is intrinsically rounder. Although not statistically significant, the Milky Way DM halo is likely to show a similar trend, as is suggested by our model with varying $p_{\rm DM}(r)$ and $q_{\rm DM}(r)$. The radial variation in the DM halo shape that we found, although tentative, is consistent with recent results from the modelling of the Sagittarius stream (Vasiliev et al., 2021).
@@ 5.2 Arrangement of the satellite system
Figure 10: Orbit configuration of satellite galaxies for the Milky Way and the four analogues highlighted in Fig.8 and Fig.9. We show orbits of the ten brightest or most massive satellites in YZYZ planes. The blue disc at the top left indicates the orientation of the stellar disc. The analogues exhibit striking similarities with the Milky Way that the DM halo coincides with a plane of satellites and vertically aligns with the Galactic disc.
@@
The so-called ‘satellite plane’ of the Milky Way is especially notable in the orbits of the most massive satellites (Sawala et al., 2023). We first illustrate the orbital configuration of the massive satellite galaxies in the Milky Way, as well as those in the 13 TNG50/Auriga galaxies whose DM shapes fall within the $3\sigma$ confidence level of the Milky Way.
@@
For the Milky Way, we selected the ten brightest satellites and used their observed 3D positions and velocities as initial conditions, as per Pace et al. (2022), to compute their orbital paths within our derived optimal gravitational potential. For each of the TNG50/Auriga galaxies, we took the ten most massive satellites and used their actual trajectories back approximately one orbital period from $z=0$ in most cases. For the satellites that have been orbited for many periods and are already very close to the galaxy centre at $z=0$, we took an orbital period from 3 Gyr ago for better visualisation.
@@
In Fig.10, we show these four galaxies that have a nearly oblate outer DM halo vertical to the stellar disc: TNG50 501208, TNG50 546474, Auriga 12, and TNG50 483594. They exhibit vertical satellite planes similar to the Milky Way; the first three are the only galaxies with DM shapes that fall within the $1\sigma$ confidence level of the Milky Way. There are no clear satellite planes in the remaining nine galaxies within $3\sigma$ (see Appendix A).
@@
In general, a correlation is evident between the DM halo shape and the arrangement of the satellite systems in TNG50 galaxies (see Appendix A). Most TNG50 galaxies have oblate DM halos with long-intermediate axes aligned with the stellar disc and paired with coplanar satellite systems. While galaxies with almost spherical DM halos display satellites distributed in a nearly isotropic fashion, those with DM halos aligned vertically to the stellar disc tend to have satellite systems organised vertically as well. This correlation is expected as a result of the DM halo assembling over time through the accretion of satellite galaxies with similar net angular momentum, consistent with previous findings from other cosmological simulations (Shao et al., 2016).
@@
5.3 A vertically orientated dark matter halo in alignment with a satellite plane results from a flip of the stellar disc
@@
Figure 11: Spiral-in minor merger flips the stellar disc illustrated by TNG50 501208. The left column show orbits of the ten most massive mergers coloured their maximum history mass. The dashed red curve with $M_{\rm max}=5.3\times 10^{10}$M  is a major merger finished at about 8.5 Gyr ago, and the solid red curve with $M_{\rm max}=7.1\times 10^{10}$M  is the most massive minor merger followed. The rest columns show the DM distribution (the grey shadow), the projection of the stellar disc (the blue ellipse with size enlarged five times for visualisation), and the orbit progress of the most massive minor merger (solid red curve) at different snapshots in the history. The axis ratios of the DM halo measured at $r=50$ kpc are labeled. A thin stellar disc was formed right after the major merger finished at about 8.5 Gyr ago. At about t=67t=6-7 Gyr ago the disc was aligned almost co-planar with the long-intermediate axis plane of the DM halo and orbital plane of the satellites. The disc was flipped in the last seven gigayears, associated with the spiral-in minor merger (solid red curve), while the general shape of DM halo (illustrated by $q_{\rm DM}$ and $p_{\rm DM}$ measured at 50 kpc in a fixed co-ordinate) and satellite arrangement remain unchanged, resulting in the stellar disc vertical to the DM halo and satellite orbit plane at present (also see Movie S1).
@@
Figure 12: Evolution history of TNG50 501208 (left) and TNG50 546474 (right). We describe the four panels with TNG50 501208. Top: Assembly history of the galaxy. The solid thick curves represent two of the most massive mergers the galaxy experienced. The dashed curves represent small mergers. TNG50 501208 has assembled $80\%$ of its DM mass and $72\%$ of its stellar mass at $t=6$ Gyr ago. Second row: Distance of the satellite to the galacetic center as a function of time. The solid black curve is a major merger ended at $t\sim 8.5$ Gyr ago. The red curve is the following most massive minor merger with $M_{\rm max}=7.1\times 10^{10}$ M. The vertical dashed red line indicates the time at which it reached the pericenter for the first time. Third row: Variation in the disc spin direction as a function of time. The disc was co-aligned with the long-intermediate axis plane of the DM halo at $t\sim 6$ Gyr ago, and it has tilted $80^{\circ}$ in the past 6 Gyr. Bottom: Evolution of the DM axis ratios, with $p_{\rm DM}=Y/X$ and $q_{\rm DM}=Z/X$. The DM shape remain largely unchanged in the past 8 Gyr. The four panels in the right show a similar evolution history of TNG50 546474.
@@
We further traced the evolutionary history of the four Auriga/TNG50 galaxies, TNG50 501208, TNG50 546474, TNG50 483594, and Auriga 12, and found that the distinct vertical alignment of the DM halo and satellite plane is due to the tilt of the stellar disc with a large angle, which we call a ‘flip of the disc’. To check the evolutionary history, we paid special attention to the merger history, the evolution of the stellar disc, and the shape of the DM halo. We calculated the spin direction of all stars at $r<20$ kpc at each snapshot in the history and defined $\theta_{\rm disc}$ as the angle between the spin vector and its projection line in the XY plane. In this definition, we have $\theta_{\rm disc,z=0}=90^{\circ}$. We then calculated the DM axis ratios at 50 kpc ($p_{\rm DM}=Y/X$, $q_{\rm DM}=Z/X$) in different snapshots. $XYZ$ is the co-ordinate defined at $z=0$, with the $Z$ axis perpendicular to the disc, as we defined previously.
@@
As is illustrated in Fig.11 and quantified in Fig. 12, the last major merger of TNG50 501208 ended $t\sim 8.5$ Gyr ago. A stellar disc was formed gradually after the major merger. At $t\sim 6-7$ Gyr ago, the stellar disc was almost aligned in the long-intermediate axis plane of the DM halo; if not perturbed by other mergers, the system would stabilise in this configuration. The galaxy was followed by a series of minor mergers, the most massive one having $M_{\rm max}=7.1\times 10^{10}$ Mcoming in on an orbit that is retrograde to the disc rotation. At $t=6$ Gyr ago, when the massive minor merger reached its pericenter for the first time, the disc started to tilt, initiating continuous tilting in the subsequent 6 Gyrs. The disc has since flipped about $80^{\circ}$ until $z=0$, with an average rate of $\sim 13^{\circ}$ per Gyr, while it was not significantly perturbed. Although the satellite perturbed the DM halo’s shape at pericenter, the DM halo quickly re-stabilised. The mass accreted from this satellite and other minor mergers continuously built the DM halo along the orbit, consistent with the pre-existing DM halo shape; the axis ratios of the DM halo measured at $r=50$ kpc remained largely unchanged. The disc flip resulted in the DM halo and stellar disc’s vertical alignment at $z=0$.
@@
The satellite galaxy that spiralled inwards brings a substantial net angular momentum. It is important to note that a tilt rate of $\sim 13^{\circ}$ per Gyr corresponds to about 2 km/s at a radius of $r=10$ kpc, indicating a very minor rigid motion of the disc. This scenario requires torques that transfer less than $\lesssim 1\%$ of the angular momentum of the satellite galaxy, which has a total mass comparable to that of the stellar disc. In our analysis of 13 TNG/Auriga galaxies within a $3\sigma$ threshold, eight galaxies exhibit significant disc tilts, with tilt angles ranging from $30^{\circ}$ to $100^{\circ}$. Seven of them are likely caused by minor mergers with $M_{\rm max}>10^{10}\,M_{\odot}$, mostly spiralled in (see appendix A). Disc flips are fairly common in cosmological simulations (Dillamore et al., 2022; Earp et al., 2017; Bett and Frenk, 2012). Direct N-body simulations show that the torque on the stellar disc from a satellite galaxy is a promising mechanism that induces the disc tilt (Dodge et al., 2023). Previous analysis of Auriga galaxies shows that disc tilt is correlated with the time duration of close satellite interactions (Gómez et al., 2017a), consistent with our results. It is also found that even low-mass satellites that penetrate the outer regions of a galaxy can significantly perturb and tilt a host galactic disc (Gómez et al., 2017b), due not only to direct tidal perturbation but also to the generation of asymmetric features in the DM halo that can be efficiently transmitted to its inner regions.
@@
In a similar pattern, the evolution of TNG50 546474 is presented in the right panel of Fig. 12. This galaxy underwent a major merger that ended around $t\sim 7.5$ Gyr ago, immediately followed by the most massive minor merger with $M_{\rm max}=1.4\times 10^{10}$M. At $t\sim 7$ Gyr ago, after the reformation of the disc post-major merger and stabilisation of the DM halo, the disc and DM halo were nearly coplanar and aligned. Subsequently, the DM halo’s shape saw minimal change, although it gained about $30\%$ of its mass in the upcoming 7 Gyr through minor mergers on orbits consistent with the pre-existing DM halo shape. At the same time, the disc has tilted approximately $60^{\circ}$ with an average rate of $\sim 8^{\circ}$ per Gyr.
@@
Similar figures for Auriga 12 and TNG50 483594 are shown in Figure 13. Auriga 12 has had more frequent recent mergers. Its stellar disc was coplanar aligned with the DM halo at $t\sim 8-10$ Gyr ago; the torques from the combination of a few subsequent minor mergers have tilted the disc by $\sim 90^{\circ}$. The DM halo was influenced but has maintained a similar shape since about 10 Gyr ago.
@@
The DM shape of TNG50 483594 is outside the $1\sigma$ but within the $3\sigma$ confidence level of the Milky Way. This galaxy is special in that it was quickly assembled with many head-on mergers at $t\gtrsim 8$ Gyr ago, but without a GSE-like major merger, which resulted in a rather spherical DM halo. Due to frequent mergers in the early Universe, it has quickly assembled stellar mass and halted star formation. With a near-spherical DM halo, its stellar disc was arbitrarily orientated at t=610t=6-10 Gyr ago. The subsequent minor mergers exhibit a satellite plane whose contribution has not significantly changed the DM halo shape until $t\sim 6$ Gyr ago, along with a massive minor merger with $M_{\rm max}=4.5\times 10^{10}$M. The minor mergers built the halo along their orbits and tilted the disc by about $30^{\circ}$ in another direction, leading to the vertical alignment of the DM halo and the disc at $z=0$. In this case, the tilt of the disc is not the only reason; it strengthens the misalignment and contributes to the vertical alignment of the stellar disc and the satellite distribution observed at $z=0$. Even without the disc tilt, there is still a misalignment of the stellar disc and the satellite plane due to variations in the filament network along with mass accreted into the galaxy, as is shown in (Shao et al., 2021).
@@
Despite the tilting of the disc, the DM halo of the inner regions ($r\lesssim 20$ kpc) was tilting with the stellar disc. As a result, the DM halos of these galaxies tend to be twisted, being oblate in the inner regions, and more vertically aligned with the disc in the outer regions, as has already been shown in Figure 9, consistent with the trend we obtained for the Milky Way.
@@
The three $1\sigma$ galaxies, TNG50 501208, TNG50 546474, and Auriga 12, would transform into a typical galaxy with an oblate DM halo by hypothetically tilting their stellar disc (and inner DM halo at $r\lesssim 20$ kpc) back to the long-intermediate axis plane of the DM halo in the outer regions. Note that TNG50 403894 also has a twisted DM halo; however, its transition radius, similar to other cases of twisted halos caused by variations in filament accretion (Shao et al., 2021), is larger than the three cases caused purely by the tilt of the disc (see Fig.9).